OpenStudy (el_arrow):

use the disk method to find the volume of the solid formed by revolving the region bounded by the graph of y=x^3, the y-axis and the lines y=1 and y=8 about the y-axis

OpenStudy (el_arrow):

i dont know how to do this problem can you help please

OpenStudy (anonymous):

Start by taking \[ dV= \pi x^2 dy \]

OpenStudy (anonymous):

\[ y= x^3\\ dy = 3 x^2 dx \]

OpenStudy (el_arrow):

okay but why a 2 its x^3

OpenStudy (anonymous):

x is the radius of the disk

OpenStudy (anonymous):

\[ dV= \pi x^2 ( 3 x^2 dx)= 3 \pi x^4 dx \]

OpenStudy (anonymous):

\[ V= \int_1^2 3 \pi x^4 dx \] Compute the integral and you are done

OpenStudy (el_arrow):

where do you get the 2 and 1? you dont use the y=1 and y=8?

OpenStudy (anonymous):

if y=1, then x=1 if \( y = x^3=8 \) then x =2

OpenStudy (el_arrow):

oh okay i see

OpenStudy (el_arrow):

thank you

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

@eliassaab where did you get that theory from? About the dV=pi*x^2*dy?

OpenStudy (anonymous):

What is the volume of a disk of radius x and thickness dy?

OpenStudy (anonymous):

@doulikepiecauseidont

OpenStudy (anonymous):

pi*x^2*dy